OpenStudy (kinggeorge):

[SOLVED] Let's see some creativity! Without using "+" make the number 9 using only three 3's, and no other digits using any mathematical symbol you want. So "3+3+3=9" and similar expressions are off limits. Here are a couple of the examples I've found so far: \[\frac{3^3}{3}\]\[\sqrt{3^3\cdot3}\]I know of several more possibilities (not including various possible applications of negation). Which ones can you get? PS: \(-(-\sqrt{3^3\cdot3})\), \(-(-3-3-3)\) and similar don't count.

OpenStudy (anonymous):

\[\sqrt {3 \times 3} \times 3\]

OpenStudy (anonymous):

\[3^0 \times 3 \times 3\]

OpenStudy (anonymous):

\[\sqrt{3!3!3!}\]

OpenStudy (kinggeorge):

@abb50 That has a 0 in it. Let me correct the problem to specify against that. @Cortegu10 \(\sqrt{3!3!3!}=\sqrt{6^3}=6\sqrt{6}\neq9\)

OpenStudy (anonymous):

oh damn i tried :)

OpenStudy (unklerhaukus):

this probably dosent count as there are four 3's best i can do \[3 \log_3(3)^3=9\]

OpenStudy (kinggeorge):

Too many 3's there :( Nice try though. I know of at least 3 more expressions using increasingly convoluted nested functions.

OpenStudy (kinggeorge):

Hint (for a couple): Keep thinking with factorials and exponents. And remember, "-" isn't completely ruled out. Just don't abuse it.

OpenStudy (experimentx):

\[ \sqrt{3!3!} + 3\]

OpenStudy (kinggeorge):

Excellent. I hadn't thought that one. Also, feel free to abuse the floor and ceiling function.

OpenStudy (experimentx):

\[ 3^{\frac{3!} 3}\]

OpenStudy (kinggeorge):

Excellent once again!

OpenStudy (experimentx):

LOL .. not sure if it works \[ \left \lfloor {3*3 + \sin 3} \right \rfloor \] \[ \left \lceil {3*3 + \cos 3} \right \rceil \]

OpenStudy (apoorvk):

greatt!!!! sin3 works!!

OpenStudy (experimentx):

perhaps log 3 too :D

OpenStudy (kinggeorge):

If you could do those without the "+" sign, those would be accepted. I'm pretty sure you can get rid of it however.

OpenStudy (apoorvk):

sin3 is something between 0 and 1, so it will. cos3 unfortunately is negative.

OpenStudy (apoorvk):

hmm.. the plus sign..

OpenStudy (kinggeorge):

Just to be clear, (Partial) list of lesser known functions I will accept: \[\lfloor9.5=9\rfloor\]\[\lceil8.5=9\rceil\]\[^33=3^{3^3}\] Also, I will accept \(\ln\) for \(\log_e\) and \(\log\) for \(\log_{10}\) as allowable functions.

OpenStudy (experimentx):

Ah great this works ceil 9^(-cos(3))

OpenStudy (experimentx):

\[ \lceil (3*3)^{-\cos 3} \rceil \]

OpenStudy (kinggeorge):

This is probably the most convoluted solution I've come up with \[\left\lceil \sqrt[3]{\left(\left(\lfloor\sqrt3\rfloor3\right)!\right)!}\right\rceil\]

OpenStudy (lgbasallote):

@experimentX \[\sqrt{3!3!} \times 3\]..youarent allowed to use + lol

OpenStudy (lgbasallote):

i mean \[\sqrt{3!3!} + 3\]

OpenStudy (lgbasallote):

you cant use plus

OpenStudy (kinggeorge):

Should've caught that :/

OpenStudy (kinggeorge):

I have at least 4 more solutions no one has posted so far =D

OpenStudy (experimentx):

\[ \left\lceil \sqrt[3]{\left( \sqrt{\left(3!3!\right)}\right)!}\right\rceil\]

OpenStudy (anonymous):

\[y=\left(\frac{x^3}{3}\right)\]dy/dx at x=3 lol but it's identical to (3^3)/3. Not sure if it counts.

OpenStudy (experimentx):

\[ (3*3)^{\lfloor \sqrt 3\rfloor }\]

OpenStudy (experimentx):

i bet sqrt 3 can be replaced with ln log using ceil ..

OpenStudy (kinggeorge):

Probably, but let's try and get things that look new, and not just replacing one part with another.

OpenStudy (kinggeorge):

I still have 3 more solutions that look different from any posted above.

OpenStudy (experimentx):

lol ... this ceil function is so useful http://www.wolframalpha.com/input/?i=3%5Eceil%28+sqrt%283*ln+3%29%5D%29

OpenStudy (anonymous):

e^(3ln(3))/3

OpenStudy (kinggeorge):

How about \[\large 3^{\lceil\sqrt3\rceil\cdot\lfloor\sqrt3\rfloor}\]

OpenStudy (experimentx):

lol ... i think we should ban usage of ceil http://www.wolframalpha.com/input/?i=3*+ceil%28log+%283*3%29%29

OpenStudy (kinggeorge):

Alright. Let's ban the ceiling function for now. What else have we got? btw, I still have 2 more different solutions

OpenStudy (experimentx):

is e allowed??

OpenStudy (kinggeorge):

Let's restrict it so we don't have \(e\), \(\pi\), \(\phi\), or other constants like that for now.

OpenStudy (kinggeorge):

Also, let's stay out of integrals and derivatives for now as well. Maybe I'll do this again with those allowed.

OpenStudy (experimentx):

this seem to have interesting result http://www.wolframalpha.com/input/?i=floor%28ln%28%283*3%29%21%29+-+3%29

OpenStudy (anonymous):

33 (mod 3)

OpenStudy (kinggeorge):

\(33\equiv0\pmod3\), although you have a case in that \(9\equiv0\pmod3\) as well.

OpenStudy (kinggeorge):

I've got to go to bed now. Keep posting solutions, and I'll post the ones I have left tomorrow.

OpenStudy (unklerhaukus):

\[\frac{3\times3!}{\Gamma(3)}=9\]

OpenStudy (experimentx):

Nice idea

OpenStudy (kinggeorge):

Here are the other ideas I've had that look different (mostly) from previous answers\[3^{3!}-(3!)!\]\[\lfloor \log(^33))\rfloor-3\]Recall that \(^33=3^{3^3}=3^{27}\)